Circular No. 8510

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 GF_{8} (LONEOS) An apparently asteroid object discovered by the LONEOS project (dis-

covery observation below) and linked by G. V. Williams, Minor Planet Center, first with apparently asteroidal Apr. 3 Siding Spring observations (and designated 2005 GF_{8} on MPS 130867) and later with Apr. 7 LINEAR and Apr. 11 Spacewatch observations, was posted on the ‘NEO Confirmation Page’ on Apr. 11. Following a request to the Spacewatch observers, J. Scotti and M. Block report that their CCD images taken on Apr. 12.32– 12.35 show the object (in the glare of a bright background star) as having a tail toward the east-southeast of the apparently stellar condensation. G. R. Jones (Tucson, AZ, 0.40-m reflector) reports that a 12′′ tail in p.a. 305^{o} is visible on his CCD images from Apr. 13.2 (noting a red magnitude for the object of 17.5). P. Birtwhistle (Great Shefford, Berkshire, England, 0.30-m reflector) writes that his CCD images taken on Apr. 11.96 and 12.08 in poor seeing show the object to be less concentrated than a star of similar brightness, with a diameter of 6′′-9′′.

2005 UT  α_{2000}  δ_{2000}  Mag.
Apr. 2.42548  14^{h}03^{m}47.^{s}65  −14^{o}19′53′′.5  17.8

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2005-G88.

T = 2005 Aug. 19.609 TT  ω = 286.^{o}248
e = 0.52072  Ω = 315.009 }2000.0
q = 2.81341 AU  i = 1.186
a = 5.87012 AU n^{o} = 0.069300 P = 14.2 years

SUPERNOVA 2005ba IN NGC 3746 R. J. Foley, M. Ganeshalingam, D. S. Wong, B. J. Swift, and A. V.

Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained Apr. 11 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2005ba (IAUC 8503) is of type II, probably within 2–3 weeks after the explosion. The spectrum consists of a blue, relatively featureless continuum with P-Cyg H lines. Adopting the NED redshift of 9022 km/s, the expansion velocity derived from the minimum of the Hβ absorption line is ∼ 11000 km/s.

2005 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green